# HyGra: Accelerating Network-State Simulation for LLM Training in DCNs via Adaptive Packet-Flow Granularity

### 13 Mar 2026 v1

@@ Abstract
In recent years, large language models (LLMs) have driven substantial intelligent transformation across diverse industries. Commercial LLM training is typically performed over data center networks (DCNs) comprising hundreds to thousands of GPUs, with multiple devices collocated per node. As network scale expands, inter-node communication becomes a primary bottleneck to training efficiency. Network-state simulators therefore play a crucial role by enabling cost-effective evaluation of network configurations and parallelization strategies through faithful emulation of DCN dynamics during LLM training. However, existing simulators are constrained by a efficiency-fidelity tradeoff, as packet-level simulators (PLSs) incur prohibitive runtime overhead, whereas flow-level simulators (FLSs) compromise essential modeling accuracy. In this paper, we develop HyGra, a hybrid-granularity network-state simulator that exploits intrinsic network dynamics in LLM training to adaptively switch simulation granularity. Specifically, HyGra employs packet-level simulation during non-steady phases with transient fluctuations and flow-level simulation during steady phases with periodic patterns, thereby accelerating execution while preserving high fidelity. Moreover, it requires no specialized hardware, supports single-machine deployment, and is compatible with existing simulators. Experiments based representative commercial LLM workloads, including ChatGPT, DeepSeek, and Qwen, show that HyGra achieves up to 15.4× speedup under single parallelization strategy and 7.8× under hybrid parallelization strategies while maintaining high accuracy.
@@ III-B Overview of HyGra
To address above challenges while ensuring compatibility with existing simulators, HyGra adopts a hierarchical architecture comprising a control layer and an execution layer as illustrated in Fig. LABEL:HyGra. This decoupled design reflects the distinct requirements of the two tasks: the control layer conducts global observation and decision making to identify granularity switching points, while the execution layer conducts network-state simulations at different granularities, as well as the transitions between them.
@@ III-B1 Parameter Input
As illustrated in Fig. LABEL:HyGra , the parameter initialization phase configures the simulation environment using network topology parameters such as the number of nodes and topology type, hardware settings such as link bandwidth and port speed, and parallelization strategies including PP, TP, and DP.
@@ III-B2 Control Layer
As shown in Fig. LABEL:HyGra , the responsibility of this layer is to establish the temporal boundaries of flow-level simulation and thus consists of two operations:
@@ III-B3 Execution Layer
As illustrated in Fig. LABEL:HyGra , once a granularity transition is triggered, the execution layer performs the network state transformations. It abstracts packet-level state into flow-level representations when entering flow granularity, and reconstructs packet-level state when resuming packet granularity. This layer comprises the following operation modules:
@@ IV Architecture of Control Layer
As the decision-making core of HyGra, the control layer determines when to switch simulation granularity. As shown in Fig. LABEL:HyGra , the control layer identifies the start and end points of steady phase, which directly determine the duration of flow-granularity simulation, thus influencing both efficiency and fidelity of the simulator as discussed in the first challenge of Section III-A. To select appropriate switching points, HyGra employs a practical steady-state recognition and a duration estimation methods, both relies solely on network bandwidth and queue-length information.
@@ VI-A5 Baselines
Packet-level Simulator (PLS): This baseline relies exclusively on the packet-granularity simulation engine (E1 in Fig. LABEL:HyGra) and serves as the upper bound on fidelity.
Flow-level Simulator (FLS): This baseline solely operates on the flow-granularity simulation engine (E2 in Fig. LABEL:HyGra). As discussed in Section II-A, it sacrifices fidelity to maximize execution speed, thus serving as a theoretical upper bound on simulation efficiency in our experiments.
@@ VI-B2 Fidelity
VI-C Performance Under Mixed Parallelization Strategies
VI-D Necessity of State Restoration
As discussed in Section V-C, flow-to-packet state restoration is crucial for preventing state loss when reverting from flow-level to packet-level simulation. Table IV quantifies its impact on fidelity using the maximum per-flow FCT error. Without state restoration, the maximum FCT error reaches 15.32% and 12.48% for DeepSeek and Qwen at 128 nodes, increasing to 26.63% and 16.98% at 512 nodes. Enabling state restoration significantly reduces these errors to 4.6% and 1.2% at 128 nodes, and to 6.7% and 2.7% at 512 nodes. This result stems from residual bandwidth fluctuations during steady phases that induce queue-depth variations. As flow-granularity simulation neglects switches and does not capture queue variations, the queue states at entry to and exit from flow-level simulation are generally inconsistent.
VI-E Comparison With Acceleration Based on Parallel Deployment

### 19 Mar 2026 v2

@@ Abstract
In recent years, large language models (LLMs) have driven substantial intelligent transformation across diverse industries. Commercial LLM training is typically performed over data center networks (DCNs) comprising hundreds to thousands of GPUs, with multiple devices collocated per node. As network scale expands, inter-node communication becomes a primary bottleneck to training efficiency. Network-state simulators therefore play a crucial role by enabling cost-effective evaluation of network configurations and parallelization strategies through faithful emulation of DCN dynamics during LLM training. However, existing simulators are constrained by a efficiency-fidelity tradeoff, as packet-level simulators (PLSs) incur prohibitive runtime overhead, whereas flow-level simulators (FLSs) compromise essential modeling accuracy. In this paper, we develop HyGra, a hybrid-granularity network-state simulator that exploits intrinsic network dynamics in LLM training to adaptively switch simulation granularity. Specifically, HyGra employs packet-level simulation during non-steady phases with transient fluctuations and flow-level simulation during steady phases with periodic patterns, thereby accelerating execution while preserving high fidelity. Moreover, it requires no specialized hardware, supports single-machine deployment, and is compatible with existing simulators. Experiments based representative commercial LLM workloads, including ChatGPT, DeepSeek, and Qwen, show that HyGra achieves up to 15.4 speedup under single parallelization strategy and 7.8 under hybrid parallelization strategies while maintaining high accuracy.
@@ III-B Overview of HyGra
To address above challenges while ensuring compatibility with existing simulators, HyGra adopts a hierarchical architecture comprising a control layer and an execution layer as illustrated in Fig. 4. This decoupled design reflects the distinct requirements of the two tasks: the control layer conducts global observation and decision making to identify granularity switching points, while the execution layer conducts network-state simulations at different granularities, as well as the transitions between them.
@@ III-B1 Parameter Input
As illustrated in Fig. 4 , the parameter initialization phase configures the simulation environment using network topology parameters such as the number of nodes and topology type, hardware settings such as link bandwidth and port speed, and parallelization strategies including PP, TP, and DP.
@@ III-B2 Control Layer
As shown in Fig. 4 , the responsibility of this layer is to establish the temporal boundaries of flow-level simulation and thus consists of two operations:
@@ III-B3 Execution Layer
As illustrated in Fig. 4 , once a granularity transition is triggered, the execution layer performs the network state transformations. It abstracts packet-level state into flow-level representations when entering flow granularity, and reconstructs packet-level state when resuming packet granularity. This layer comprises the following operation modules:
@@ IV Architecture of Control Layer
As the decision-making core of HyGra, the control layer determines when to switch simulation granularity. As shown in Fig. 4 , the control layer identifies the start and end points of steady phase, which directly determine the duration of flow-granularity simulation, thus influencing both efficiency and fidelity of the simulator as discussed in the first challenge of Section III-A. To select appropriate switching points, HyGra employs a practical steady-state recognition and a duration estimation methods, both relies solely on network bandwidth and queue-length information.
@@ VI-A5 Baselines
Packet-level Simulator (PLS): This baseline relies exclusively on the packet-granularity simulation engine (E1 in Fig. 4) and serves as the upper bound on fidelity.
Flow-level Simulator (FLS): This baseline solely operates on the flow-granularity simulation engine (E2 in Fig. 4). As discussed in Section II-A, it sacrifices fidelity to maximize execution speed, thus serving as a theoretical upper bound on simulation efficiency in our experiments.
@@ VI-B2 Fidelity
VI-C Performance Under Mixed Parallelization Strategies
VI-D Necessity of State Restoration
As discussed in Section V-C, flow-to-packet state restoration is crucial for preventing state loss when reverting from flow-level to packet-level simulation. Table V quantifies its impact on fidelity using the maximum per-flow FCT error. Without state restoration, the maximum FCT error reaches 15.32% and 12.48% for DeepSeek and Qwen at 128 nodes, increasing to 26.63% and 16.98% at 512 nodes. Enabling state restoration significantly reduces these errors to 4.6% and 1.2% at 128 nodes, and to 6.7% and 2.7% at 512 nodes. This result stems from residual bandwidth fluctuations during steady phases that induce queue-depth variations. As flow-granularity simulation neglects switches and does not capture queue variations, the queue states at entry to and exit from flow-level simulation are generally inconsistent.
VI-E Comparison With Acceleration Based on Parallel Deployment